OpenStudy (anonymous):

quick help with another homogeneous linear equation with initial conditions y''+y'=0 y(0)=2, y'(0)=1

OpenStudy (anonymous):

yeah this one is throwing me off :/ i think its \[r^2+r=0\]

OpenStudy (anonymous):

@satellite73 when you aren't busy please!

OpenStudy (anonymous):

@DoYourHomework well i didn't get very far into it without trouble again :/

OpenStudy (anonymous):

or should it be r^2+r+0=0? not sure if that makes a difference...it only would with quadratic equation

OpenStudy (anonymous):

these seem so easy in class and when i do homework it kicks my A**

OpenStudy (anonymous):

roots are r = -1 and r = 0

OpenStudy (anonymous):

So \[y = c_{1}e^{-x}+c _{2}\]

OpenStudy (anonymous):

\[y' = -c_{1}e ^{-x}\]

OpenStudy (anonymous):

so C1=-1?

OpenStudy (anonymous):

\[2 = c1+c2\] \[1= -c1\]

OpenStudy (anonymous):

yes :D

OpenStudy (anonymous):

ok it was just the r^2+r=0 that threw me off :/ thanks for your help again @DoYourHomework !

OpenStudy (anonymous):

np!